# Proprioceptive State Estimation for Amphibious Tactile Sensing

### v1

@@ Abstract
Keywords Soft Robotics  $\cdot$ Tactile Sensing  $\cdot$ State Estimation
@@ 1 Introduction
For robots to interact with the physical environment, the moment of touch holds the truth of dynamic interactions [1]. For most living organisms, the skin plays a crucial role in translating material properties, object physics, and interactive dynamics through sensory receptors into chemical signals processed by the brain [2] to formulate a feeling of the external environment (exteroception) [3] and the bodily self (proprioception) [4]. One stream of research focuses on developing novel tactile sensors [5], aiming at replicating the skin’s basic functionality with comparable or superior performances [6]. Another stream of research considers robots while developing or utilizing tactile sensors [7], where integrating with robotic systems becomes a design challenge that requires an interdisciplinary approach [8], fostering a growing interest in tactile robotics among academia and industry [9].
@@ 1.1 Towards Dense Sensing for Tactile Robotics
Tactile sensory generally involves many properties that can be digitized for robotics [10]. For mechanics-based dynamics and control, the interactive force and torque on the contact surface are a primary concern in robotics [11], which could be generally approached as 1) Point-wise Sensing for 6D forces and torques (FT), 2) Bio-inspired Sparse Sensing Array and 3) Visuo-Tactile Dense Image Sensing, where a certain level of material softness or structural geometry are leveraged for an enhanced representation of the mechanic interactions into tactile data.
@@
Point-wise Sensing for 6D FT: Estimating forces at contact points is paramount in robotic systems, enabling awareness of physical interaction between the robot and its surrounding objects [12]. Robotic research, especially when dynamics and mechanics are involved, is generally more interested in utilizing the force-and-torque properties for manipulation problems by robotic hands [13] or locomotion tasks by legged robots [14]. For a single reference point, the FT properties could be succinctly represented by a 6D vector of the forces along and the torques about the three orthogonal axes, making it comparable to the joint torque sensing in articulated robotic structures. However, the shortcut between physical contact and a point-wise 6D FT measurement may not capture the full extent of contact information for further algorithmic tasks [15].
@@
Bio-inspired Sparse Sensing Array: Similar to the biological skin’s super-resolutive mechanoreception for tactile sensing, a common approach in engineering is to place an array of sensing units on the interactive surface [16]. Instead of going for a localized 6D force and torque contact information, researchers usually tackle the problem with enhanced pressure sensing across its entire surface from spatially distributed sensing elements [17]. As a result, one can build models or implement learning algorithms to achieve super-resolution by sampling the discrete sensory inputs, producing a continuous estimation of the tactile interaction on the surface at a much higher resolution than the sensing array arrangement. Recent research [18] shows that one can leverage magnetic properties to achieve de-coupled normal and shear forces and simultaneously achieve super-resolutive tactile sensing of the normal and frictional forces for high-performing robotic grasping.
@@
Visuo-Tactile Dense Image Sensing: Vision-based tactile sensing recently emerged as a popular approach to significantly increase the sensing resolution by leveraging the modern imaging process to visually track the deformation of a soft medium as the interface of physical interaction [19, 20, 21], eliminating the need for biologically inspired super-resolution [22]. Robotic vision has already become a primary sensing modality for advanced robots with the maturity of modern imaging technologies within a compact form factor [23] accessible to various software and algorithm libraries for real-time processing. While the high resolution of modern cameras offers significant advantages, the infinite number of potential configurations of the soft medium introduces a significant challenge [24].
@@ 1.2 Proprioceptive State Estimation
For tactile applications in robotics, proprioceptive perception of the joint position and movement of the body plays a critical role in achieving state estimation. The tactile interface is a physical separation between the intrinsic proprioception concerning the robot and the extrinsic perception concerning the object-centric environment. For the following, we focus on vision-based tactile sensing, which also applies to analyzing the other methods reviewed above.
Intrinsic Proprioception in Tactile Robotics: For vision-based intrinsic proprioception, the analysis is usually centered on estimating the state of the soft medium during contact, inferring tactile interaction [25]. To establish a physical correspondence between a finite parameterization state estimation model and an infinite configuration of soft deformation [26], markers that are easy to track are often used to discretize the displacement field of soft mediums. In [27], a simple blob detection method is introduced to track uniform distributed markers in a planar transparent soft layer for deformation approximation. Advanced image analysis [28] is also adopted to utilize machine learning algorithms to extract high-level deformation patterns from randomly spread markers over the entire three-dimensional volume of soft medium for robust state estimation [29]. Recent research [30] shows a promising approach to integrate physics-based models that capture the dynamic behavior of the soft medium under deformation.
Extrinsic Perception for Tactile Robotics: For extrinsic perception, the focus is shifted to estimating the object-level information, including object localization, shape and dynamics parameters, etc. [31], where tactile sensing data could be used for task-based manipulation and locomotion. Using contact to estimate an object’s global geometry is instrumental for intelligent agents to make better decisions during object manipulation [32]. Usually, tactile sensing is employed for estimating the object’s shape in visually occluded regions, thus playing a complementary role to vision sensors [33, 34]. However, in scenarios where a structured environment with reliable external cameras is unavailable or impractical, such as during exploration tasks in unstructured environments, tactile sensing can provide valuable feedback to achieve environmental awareness [35].
@@ 1.3 Amphibious Tactile Robotics
Amphibious environments present a unique and dynamic challenge for robotic systems. Robots operating in these environments must contend with vastly different physical properties, including changes in buoyancy, friction, and fluid dynamics [36]. Furthermore, the transition between water and air requires robots to adapt their sensory systems and control strategies to function effectively in each medium [37].
@@
Developing effective tactile sensors for amphibious robots presents several challenges. Sensors must be robust enough to withstand the harsh aquatic environment and be sensitive enough to detect subtle changes in water and air [38]. The transition between these two media can also cause sensor drift and require calibration to maintain accuracy [39]. Despite these challenges, there are exciting opportunities in amphibious tactile robotics, with improved sensitivity, durability, and resistance to environmental factors [40]. However, a research gap remains in developing an effective tactile sensing method with an integrated design on the finger surface or fingertip that directly applies to amphibious applications.
1.4 Proposed Method and Contributions
This paper introduces a novel vision-based proprioception approach for a soft robotic finger, enabling the estimation and reconstruction of tactile interactions in terrestrial and aquatic environments. This is achieved by leveraging the finger’s metamaterial structure, which allows for omni-directional passive adaptation, protecting delicate objects during grasping across diverse scenarios. The proposed system employs a compact in-finger camera to capture high-framerate images of the soft finger’s morphology during contact, extracting crucial tactile data. The metamaterial structure of the finger allows for omni-directional passive adaptation, safeguarding fragile objects during grasping in both water and air. The in-finger camera enables real-time extraction of tactile data from the finger’s deformation, providing valuable insights into the physical interaction. This combination of passive adaptation and real-time tactile sensing makes the proposed system well-suited for amphibious robotic grasping applications. Here are the contributions of this study:
  • Proposed and validated a model-based synthesis of morphological deformation for an omni-adaptive soft robotic finger using in-finger vision.
  • Formalized a contact-based object surface estimation approach using a spatial discretized elastic model.
  • Evaluated the proprioceptive state estimation performance of the soft omni-adaptive finger with demonstrated performance for underwater object shape estimation.
This paper is organized as follows. Section 2 provides the design of an omni-adaptive soft finger. Section 3 presents a proprioceptive shape modeling and estimation method. Section 4 addresses touch sensing for contact object shape estimation. Experiment results are shown in Section 5. In Section 6 we discuss sensorization and morphological design of the soft finger for touch sensing. Section 7 includes conclusion, limitations, and future works.
2 Omni-adaptive Metamaterial Design with In-Finger Vision
Soft grippers can achieve diverse and robust grasping behaviors with a relatively simple control strategy [41]. This study adopts a vision-based approach to enable tactile perception for a soft robotic finger, as shown in Fig. 1A. The finger geometry features a shrinking cross-sectional network towards the tip based on our previous work [42, 43], fixed on a transparent support frame by a mounting plate. A rigid plate pasted with an ArUco tag is mechanically fixed with the four lower crossbeams of the soft finger. To capture the pose of the tag, a monocular RGB camera with a field of view (FOV) of ${}^{\circ}$ is installed inside the support frame facing towards the cavity of the finger, which is capable of video recording at a high frame rate at 120 FPS (frames per second) and a resolution of 640 × 480 pixels. During the interaction between the finger and the external environment, the real-time image captured by the camera provides the pose data of the ArUco tag as rigid-soft kinematics coupling constraints for the proprioceptive state estimation of the finger.
@@ 2 Omni-adaptive Metamaterial Design with In-Finger Vision
The finger is fabricated by vacuum molding using Hei-cast 8400, a three-component polyurethane elastomer. Based on the previous experience [44], we mixed the three components with a ratio of 1:1:0, producing a hardness of 90 (Type A) to achieve reliable bending and enveloping capabilities. The proposed soft finger exhibits complex adaptive deformations, such as forward push, oblique push, and twist, as shown in Fig. 1B. Due to the capability of omni-directional adaptation on the finger surface to unknown object geometries, it enables a passive form-closure for robotic grasping [45].
3 Shape Deformation Modelling and Estimation
For the volumetric soft finger, we consider a 3D space $\mathbf{\Omega}\in{\mathbb{R}^{3}}$ filled with homogeneous elastic material. Under different boundary conditions, the distribution of internal elastic energy within the soft body varies significantly. The objective of proposed proprioceptive state estimation methodology is to accurately determine the smooth deformation map, $\Phi:\mathbf{\Omega}\rightarrow{\tilde{\mathbf{\Omega}}}$, that facilitates the geometric transformation of the soft body from its initial configuration, represented by $\mathbf{\Omega}$, to a deformed configuration, denoted as $\tilde{\mathbf{\Omega}}$. This transformation is characterized by minimizing a form of variational energy measuring the distortion of the soft body [46]. As a result, the proposed computational approach to soft finger shape reconstruction highly depends on finite element discretization and the choice of energy function that characterizes deformation.
3.1 Volumetric Parameterization of Deformation
Denote the tetrahedral mesh of discretized soft body by =$\mathcal{M}=\{\mathcal{V},\mathcal{T}\}$ where $\mathcal{V}=\{\mathbf{x}_{1},...,\mathbf{x}_{n}\}$ is the set of vertices $\mathbf{x}_{i}\in{\mathbb{R}^{3}}$, and $\mathcal{T}=\{t_{1},...,t_{m}\}$ is the set of tetrahedra elements, as shown in Fig. 2A(i).
@@ 3.1 Volumetric Parameterization of Deformation
When the soft body deforms, a collection of chosen linear approximated local deformation maps are applied to $\mathcal{M}$, over each tetrahedron element $t_{j}$ via an affine transformation:
@@
where $\mathbf{X}\in{\mathbb{R}^{3}}$ stands for all points inside element $t_{j}$, $\mathbf{A}_{t_{j}}\in{\mathbb{R}^{3\times{3}}}$ is the differential part of the deformation map and $\mathbf{b}_{t_{j}}\in{\mathbb{R}^{3}}$ is the translation part. We choose this piecewise linear deformation map over each tetrahedron for computational efficiency. Besides, high-order deformation functions can be used for better approximation [47].
@@
[…]
where $\mathbf{x}_{t_{j}}^{i}\in{\mathbb{R}^{3}}$, $i\in{\{1,2,3,4\}}$, are the deformed vertices location of $t_{j}$ tetrahedron, $\mathbf{X}_{t_{j}}^{i}\in{\mathbb{R}^{3}}$ are corresponding initial vertices location.
Given the locations of initial vertices, the deformation gradient $\mathbf{A}_{t_{j}}$ in the chosen piecewise linear transformation Eq. 1 can be expressed as a linear combination of unknown deformed element vertices location $\mathbf{x}_{t_{j}}$ as :
@@
For the discretized tetrahedral mesh $\mathcal{M}$, the collection of deformation maps $\{\Phi_{t_{j}}\}_{t_{j}\in{\mathcal{T}}}$ for all tetrahedra elements should uniquely determine the deformed shape of the soft body [48].
3.2 Geometry-Related Deformation Energy Function
To mimic the physical deformation behavior, the specific energy function form of the deformation map $\Psi(\Phi_{t_{j}})$ needs to be prescribed. Several proper forms of geometry-related deformation energies have been proposed, such as as-rigid-as-possible [49], conformal distortion [50], and isometric distortion [51].
Instead of deriving the energy of the system explicitly using constitutive relation and balance equations [52], we directly choose a symmetric Dirichlet form of energy function to characterize the deformation [53], which indicates isometric distortion and behaves well in the case of our soft finger. As the deformation should be irrelevant to the translation, the discrete element energy functional only takes augment of the gradient of each deformation maps $\{\Phi_{t_{j}}\}_{t_{j}\in{\mathcal{T}}}$ as:

### v2

@@ Abstract
Keywords Soft Robotics  $\cdot$ Vision-based Tactile Sensing  $\cdot$ State Estimation  $\cdot$ Shape Reconstruction  $\cdot$ Proprioception
@@ 1 Introduction
Proprioceptive State Estimation (PropSE) refers to the process of determining the internal state or position of a robot or a robotic component (such as a limb or joint) by measuring the robot’s internal properties [1, 2]. PropSE is particularly important in soft robotics, especially in terrestrial and aquatic environments, where the flexible and deformable nature of these robots makes traditional position and orientation sensing challenging [3]. During the robot’s physical exchange with the external environment, the moment of touch holds the truth of the dynamic interactions [4]. For most living organisms, the skin is crucial in translating material properties, object physics, and interactive dynamics via the sensory receptors into chemical signals [5]. When processed by the brain, they collectively formulate a feeling of the external environment (exteroception) [6] and the bodily self (proprioception) [7]. Towards tactile robotics, one stream of research aims at replicating the skin’s basic functionality with comparable or superior performances [8]. For example, developing novel tactile sensors [9] represents a significant research focus. Another stream of research considers robots while developing or utilizing tactile sensors [10]. It requires an interdisciplinary approach to resolve the design challenge involved [11], fostering a growing interest in tactile robotics among academia and industry [12].
We previously conducted a preliminary investigation on Vision-Based Tactile Sensing (VBTS) [13], which leverages the visual features of a series of soft metamaterial structures’ large-scale, omni-directional adaptative deformation. The design of these metamaterial structures was subsequently generalized as a class of Soft Polyhedral Networks (SPN) [14], for which high-performance proprioceptive learning in object manipulation was achieved via a node-based representation. Recent literature shows the growing adoption of volumetric representation with finite element modeling as the de facto ground truth for soft, dynamic interactions [15]. Yet, the high computational cost limits its application in robotic tasks, where real-time perception is critical [16]. Aquatic machine vision remains difficult [17] for unstructured underwater exploration with changing turbidity (relative clarity of a liquid measured by Nephelometric Turbidity Unit, or NTU). Finger-based PropSE complements aquatic machine vision by providing localized tactile perception in Simultaneous Localization and Mapping (SLAM) [18]. It is a research gap to investigate the design and learning trade-off between high-fidelity proprioceptive state estimation and real-time perception in an amphibious environment [15, 19, 20]. In such scenarios, in-finger vision with soft robotic fingers may provide a promising solution to advance the field of tactile robotics.
This paper introduces a Vision-Based Tactile Sensing (VBTS) approach for real-time and high-fidelity Proprioceptive State Estimation (PropSE) with demonstrated amphibious applications in the lab and field. This is achieved using the Soft Polyhedral Network structure with marker-based in-finger vision as the soft robotic fingers for large-scale, omni-directional adaptations with amphibious tactile sensing capability. We proposed a model-based approach for PropSE by introducing rigidity-aware Aggregated Multi-Handle (AMH) constraints to optimize a volumetric parameterization of the soft robotic finger’s morphological deformation. This enabled us to restructure the VBTS problem as an implicit surface model using Gaussian Processes for object shape reconstruction. We benchmarked our proposed method in shape reconstruction against existing solutions with verified superior performances. We also conducted experiments using commercial-grade motion-capture systems and touch-haptic devices, demonstrating our solution’s large-scale reconstruction and touch-point estimation performances. Finally, we demonstrated the application of our proposed solutions for amphibious tactile sensing in three experiments, including a shape reconstruction experiment, a turbidity benchmarking experiment, and a tactile grasping experiment on an underwater Remotely Operated Vehicle (ROV). The following are the contributions of this study:
  • Modelled Proprioceptive State Estimation (PropSE) via rigidity-aware Aggregated Multi-Handle constraints.
  • Formulated Vision-Based Tactile Sensing (VBTS) via an Implicit Surface model for object shape reconstruction.
  • Achieved PropSE for VBTS using Soft Polyhedral Networks with in-finger vision as robotic tactile fingertips.
  • Benchmarked PropSE for amphibious tactile reconstruction with demonstrated applications & testing.
This paper is organized as follows. Section 2 briefly reviews related literature about the role of proprioceptive state estimation in tactile robotics and its application in amphibious tactile sensing. Section 3 introduces the soft robotic fingertips for this study and presents our proposed model for proprioceptive state estimation via rigidity-aware Aggregated Multi-Handle constraints. This section also formulates our proposed vision-based tactile sensing method via implicit surface modeling. All experimental results are presented in Section 4, including those for benchmarking our proposed method’s performance and those conducted explicitly for amphibious tactile sensing underwater. Conclusion, limitations, and future work are enclosed in the final section, which ends this paper.
2 Literature Review
@@ 2.1 Towards Dense Sensing for Tactile Robotics
Tactile sensory generally involves many properties that can be digitized for robotics [21]. For mechanics-based dynamics and control, the interactive forces and torques on the contact surface are a primary concern in robotics [22]. It usually involves a certain level of material softness or structural deformation for an enhanced representation of the mechanic interactions as tactile data. The following are the three general research streams in this field.
2.1.1 Point-wise Sensing in 6D FT
@@
Estimating forces at contact points is paramount in robotic systems, enabling awareness of physical interaction between the robot and its surrounding objects [23]. Robotic research, especially when dynamics and mechanics are involved, is generally more interested in utilizing the force-and-torque (FT) properties for manipulation problems by robotic hands [24] or locomotion tasks by legged systems [25]. The FT properties could be succinctly represented by a 6D vector of forces and torques for a single reference point, making it comparable to the joint torque sensing in articulated robotic structures. However, the shortcut between physical contact and a point-wise 6D FT measurement may not capture the full extent of contact information for further algorithmic processing [26].
2.1.2 Bio-inspired Sparse Sensing Array
@@
Similar to the biological skin’s super-resolutive mechanoreception for tactile sensing [27], a common approach in engineering is to place an array of sensing units on the interactive surface [28]. Instead of going for a localized 6D force and torque contact information, researchers usually tackle the problem with enhanced pressure sensing across its entire surface from spatially distributed sensing elements [29]. As a result, one can build models or implement learning algorithms to achieve super-resolution by sampling the discrete sensory inputs. This approach continuously estimates the tactile interaction on the surface at a much higher resolution than the sensing array arrangement. Recent research [30] shows that one can leverage magnetic properties to achieve de-coupled normal and shear forces with simultaneous super-resolution in tactile sensing of the normal and frictional forces for high-performing grasping.
2.1.3 Visuo-Tactile Dense Image Sensing
@@
Vision-based tactile sensing recently emerged as a popular approach to significantly increase the sensing resolution [31]. This approach leverages the modern imaging process to visually track the deformation of a soft medium as the interface of physical interaction [32, 33], eliminating the need for biologically inspired super-resolution [34]. Robotic vision has already become a primary sensing modality for advanced robots [35]. The maturity of modern imaging technologies drives the hardware to be more compact while the software is more accessible to various algorithm libraries for real-time processing. While the high resolution of modern cameras offers significant advantages, the infinite number of potential configurations of the soft medium introduces a considerable challenge [36].
@@ 2.2 Proprioceptive State Estimation
For tactile applications in robotics, proprioceptive perception of joint position and body movement plays a critical role in achieving state estimation. The tactile interface is a physical separation between the intrinsic proprioception concerning the robot and the extrinsic perception concerning the object-centric environment. We focus on vision-based proprioception, which also applies to analyzing the abovementioned methods.
2.2.1 Intrinsic Proprioception in Tactile Robotics
For vision-based intrinsic proprioception, the analysis is usually centered on estimating the state of the soft medium during contact, inferring tactile interaction [37]. To establish a physical correspondence between a finite parameterization state estimation model and an infinite configuration of soft deformation [38], markers that are easy to track are often used to discretize the displacement field of soft mediums. In [39], a simple blob detection method is introduced to track uniform distributed markers in a planar transparent soft layer for deformation approximation. Advanced image analysis [40] is also adopted to utilize machine learning algorithms to extract high-level deformation patterns from randomly spread markers over the entire three-dimensional volume of soft medium for robust state estimation [41]. Recent research [42] shows a promising approach to integrate physics-based models that capture the dynamic behavior of the soft medium under deformation.
2.2.2 Extrinsic Perception for Tactile Robotics
For extrinsic perception, the focus is shifted to estimating the object-level information. Tactile sensing data such as object localization, shape, and dynamics parameters could be used for task-based manipulation and locomotion [21]. Using contact to estimate an object’s global geometry is instrumental for intelligent agents to make better decisions during object manipulation [43]. Usually, tactile sensing is employed for estimating the object’s shape in visually occluded regions, thus playing a complementary role to vision sensors [44, 45]. However, in scenarios where a structured environment with reliable external cameras is unavailable or impractical, such as during exploration tasks in unstructured environments, tactile sensing can provide valuable feedback to achieve environmental awareness [46].
@@ 2.3 Amphibious Tactile Robotics
Amphibious environments present a unique and dynamic challenge for robotic systems [47]. Robots operating in these environments must contend with vastly different physical properties, including changes in buoyancy, friction, and fluid dynamics [48]. Furthermore, the transition between water and air requires robots to adapt their sensory systems and control strategies to function effectively in each medium [49].
@@
Developing effective tactile sensors for amphibious robots presents several challenges. Sensors must be robust enough to withstand the harsh aquatic environment and be sensitive enough to detect subtle changes in water and air [50]. The transition between these two media can also cause sensor drift and require calibration to maintain accuracy [51]. Despite these challenges, there are exciting opportunities in amphibious tactile robotics, with improved sensitivity, durability, and resistance to environmental factors [52]. However, a research gap remains in developing an effective tactile sensing method with an integrated finger-based design that directly applies to amphibious applications.
3 Materials & Methods
3.1 Soft Polyhedral Network with In-Finger Vision
Soft grippers can achieve diverse and robust grasping behaviors with a relatively simple control strategy [53]. In this study, we adopted our previous work in a class of Soft Polyhedral Networks with in-finger vision as the soft robotic finger [13, 14]. As shown in Fig. 1A, the specific design is modified using an enhanced mounting plate to fix the soft finger and made waterproof for amphibious tactile sensing. The soft finger features a shrinking cross-sectional network design towards the tip, capable of omni-directional adaptation during physical interactions, as shown in Fig. 1B. We fabricated the finger by vacuum molding using Hei-cast 8400, a three-component polyurethane elastomer. Based on our previous work, we mixed the three components with a ratio of 1:1:0, producing a hardness of 90 (Type A) to achieve reliable spatial adaptation for grasping.
@@ 3.1 Soft Polyhedral Network with In-Finger Vision
An ArUco [^1] http://sourceforge.net/projects/aruco/ tag [54] is attached to the bottom side of a rigid plate mechanically fixed with the four lower crossbeams of the soft finger. A monocular RGB camera with a field of view (FOV) of 130 is fixed at the bottom inside a transparent support frame as in-finger vision, video-recording in a high frame rate of 120 FPS (frames per second) at 640 × 480 pixels resolution. When the soft robotic finger interacts with the external environment, live video streams captured by the in-finger vision camera provide real-time pose data of the ArUco tag as rigid-soft kinematics coupling constraints for the PropSE of the soft robotic finger. This marker-based in-finger vision design is equivalent to a miniature motion capture system, efficiently converting the soft robotic finger’s spatial deformation into real-time 6D pose data.
3.2 Volumetric Modeling of Soft Deformation for PropSE
Our proposed solution begins by formulating a volumetric model of the soft robotic finger in a 3D space $\mathbf{\Omega}\in{\mathbb{R}^{3}}$ filled with homogeneous elastic material. The distribution of the internal elastic energy within the volumetric elements varies significantly depending on the boundary conditions defined. The PropSE process requires an accurate determination of a smooth deformation map, $\Phi:\mathbf{\Omega}\rightarrow{\tilde{\mathbf{\Omega}}}$, that facilitates the geometric transformation of the soft body from its initial state, represented by $\mathbf{\Omega}$, to a deformed state, denoted as $\tilde{\mathbf{\Omega}}$. This transformation is characterized by minimizing a form of variational energy measuring the distortion of the soft body [55]. As a result, the PropSE performance depends on finite element discretization and the choice of energy function that characterizes deformation.
3.2.1 Volumetric Parameterization of Whole-body Deformation
We denote a tetrahedral mesh of the discretized soft body using =$\mathcal{M}=\{\mathcal{V},\mathcal{T}\}$, where $\mathcal{V}=\{\mathbf{x}_{1},...,\mathbf{x}_{n}\}$ is the set of vertices $\mathbf{x}_{i}\in{\mathbb{R}^{3}}$, and $\mathcal{T}=\{t_{1},...,t_{m}\}$ is the set of tetrahedra elements, as shown in Fig. 2A(i).
@@ 3.2.1 Volumetric Parameterization of Whole-body Deformation
When the soft body deforms, a collection of chosen linear approximated local deformation maps are applied to $\mathcal{M}$ over each tetrahedron element $t_{j}$ via an affine transformation:
@@
where $\mathbf{X}\in{\mathbb{R}^{3}}$ stands for all points inside element $t_{j}$, $\mathbf{A}_{t_{j}}\in{\mathbb{R}^{3\times{3}}}$ is the differential part of the deformation map, and $\mathbf{b}_{t_{j}}\in{\mathbb{R}^{3}}$ is the translational part. We choose this piecewise linear deformation map for computational efficiency. High-order deformation functions can be used for better approximation if needed [56].
@@
$\begin{bmatrix}\!\mathbf{A}_{t_{j}}&\!\!\!\!\mathbf{b}_{t_{j}}\!\end{bmatrix}\!\!\cdot\!\!{\begin{bmatrix}\!\mathbf{X}_{t_{j}}^{1}&\!\!\!\!\mathbf{X}_{t_{j}}^{2}&\!\!\!\!\mathbf{X}_{t_{j}}^{3}&\!\!\!\!\mathbf{X}_{t_{j}}^{4}\!\\ \!\mathbf{1}&\!\!\!\!\mathbf{1}&\!\!\!\!\mathbf{1}&\!\!\!\!\mathbf{1}\!\end{bmatrix}}=\begin{bmatrix}\mathbf{x}_{t_{j}}^{1}&\!\!\!\!\mathbf{x}_{t_{j}}^{2}&\!\!\!\!\mathbf{x}_{t_{j}}^{3}&\!\!\!\!\mathbf{x}_{t_{j}}^{4}\!\end{bmatrix},$ (2)
where $\mathbf{x}_{t_{j}}^{i}\in{\mathbb{R}^{3}}$, $i\in{\{1,2,3,4\}}$ are the deformed vertices location of $t_{j}$ tetrahedron, and $\mathbf{X}_{t_{j}}^{i}\in{\mathbb{R}^{3}}$ are the corresponding initial vertices location.
Therefore, the deformation gradient $\mathbf{A}_{t_{j}}$ in the chosen piecewise linear transformation in Eq. (1) can be expressed as a linear combination of unknown deformed element vertices location $\mathbf{x}_{t_{j}}$ using the following formulation:
@@
For a discretized tetrahedral mesh $\mathcal{M}$, the collection of deformation maps $\{\Phi_{t_{j}}\}_{t_{j}\in{\mathcal{T}}}$ for all tetrahedra elements should uniquely determine the deformed shape of the soft body [57].
3.2.2 Geometry-Related Deformation Energy Function
To mimic the physical deformation behavior, the specific energy function form of the deformation map $\Psi(\Phi_{t_{j}})$ needs to be specified. Several formulations of geometry-related deformation energies, such as As-Rigid-As-Possible (ARAP) [58], conformal distortion [59], and isometric distortion [60], have been proposed in recent literature.
Instead of deriving the energy of the system explicitly using constitutive relation and balance equations [61], we choose a symmetric Dirichlet form of energy function [62] to characterize the deformation, which indicates isometric distortion and behaves well in the case of our soft finger. Since the deformation should be irrelevant to the translation, the discrete element energy function only takes the gradient augment of each deformation maps $\{\Phi_{t_{j}}\}_{t_{j}\in{\mathcal{T}}}$ as: